[Brahms-l] Shift report 20050620 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Jun 20 2005 - 00:02:27 EDT
Shifter leader was SB
Shifters were 


1. New store since 7:30 PM. Still running for physics. Dump scheduled
for 2:15 AM<br>

2. Runs 15675 - 15681  <br>

Trig2(954k) ,Trig4(102k) ,Trig5(879k)<br>

3. TPM1 and TPM2 are off due to drift voltage problems of TPM2<br>
------------------------------------------------------<br>
Time            Blue Pol.(%)         Yellow Pol.(%)<br>
------------------------------------------------------<br>
At injection  -42.62 +/- 2.38         -44.81 +/- 2.24<br>
22:15         -41.32 +/- 2.41         -45.34 +/- 2.30<br>


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jun 20 00:02:43 2005

This archive was generated by hypermail 2.1.8 : Mon Jun 20 2005 - 00:02:55 EDT